OpenStudy (anonymous):

Leave answer in exact form unless otherwise indicated. Find the exact circumference of a circle with diameter equal to 8 ft.

OpenStudy (mathstudent55):

C = (pi)d Substitute 8 ft for d in the formula.

OpenStudy (anonymous):

never mind got it

OpenStudy (mathstudent55):

ok

OpenStudy (anonymous):

ok, I can't figure it out. either im missing a step or somthing. I do 8*3.14*2 right?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

The exact answer means you leave pi as pi and just show the answer in terms of pi. Since the formula is C = (pi)d, and d = 8 ft, C = (pi) * 8 ft = 8(pi) ft

OpenStudy (mathstudent55):

Using pi = 3.14 is an approximation. Since the problem asks for an exact answer, you must include pi in the answer.

OpenStudy (anonymous):

25.12? so just 8 times 3.14?

OpenStudy (mathstudent55):